# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ 4.1. Choice of parameters
In the sequel, additional parameters will be indispensable and their value has to be carefully chosen in order to respect all the compatibility conditions appearing in the estimations below. First, for a fixed integer $b\geqslant 7$, we let $\beta,\gamma\in(0,1)$ be small parameters satisfying
@@
which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and choosing $\beta$ small such that $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$. The last free parameter $a\in 2^{36b^{2}\mathbb{N}}$ is chosen sufficiently large such that $a^{\frac{b\gamma}{8}}\in\mathbb{N}$ and $2<a^{(\frac{b}{4}-1)b^{2}\gamma}$. The choice of $a$ ensures $f(q)\in\mathbb{N}$ and (3.6) holds. The above requirements for parameters will be used in the following sections. In the sequel, we increase $a$ in order to absorb various implicit and universal constants.
@@ 4.2. Mollification
We replace $v_{q}$ by a mollified velocity field $v_{\ell}$, we choose a small parameter $\ell\in(0,1)$ as $\ell=\lambda_{q}^{-2}$. Let $\{\phi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers on $\mathbb{R}^{3}$, and $\{\varphi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers with support in $(0,1)$. The one-sided mollifiers here is used in order to preserve adaptedness. We define a mollification of $v_{q}$ and $\mathring{R}_{q}$ in space and time by
$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
where $\phi_{\ell}=\frac{1}{\ell^{3}}\phi(\frac{\cdot}{\ell})$ and $\varphi_{\ell}=\frac{1}{\ell}\varphi(\frac{\cdot}{\ell})$. By definition it is easy to know that $z_{\ell}$ is adapted. Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ satisfies
$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{com}}),$ (4.1)
@@
where
$R_{\textrm{com}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
@@
Using the definition of $z_{q}$ (3.5) and the pathwise inductive assumption (3.8a), we can obtain the following pathwise estimate for $R_{\textrm{com}}$
$\|R_{\textrm{com}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (4.3)
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Using a mollification estimate and inductive assumption (3.8a) and (3.8c), for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$, we have
$\|R_{\textrm{com}}(t)\|_{L^{\infty}}$
Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
$|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (4.4)
$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}$
$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $a$ was chosen sufficiently large to absorb the constant. By (4.4), we also obtain
$\qquad|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (4.5)
Similarly, by inductive assumption (3.8c) and the same requirements for parameters as above, we also obtain for any $t\in\mathbb{R}$
$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.6)
@@ 4.3. Perturbation of the velocity
As usual in convex integration schemes, in order to guarantee smoothness throughout the construction, the new velocity field $v_{q+1}$ will be constructed as a perturbation of $v_{\ell}$:
$v_{q+1}=:v_{\ell}+w_{q+1},$
where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed construction with introducing transport equations and time cutoffs.
@@ 4.3.1. Flow maps and cutoffs
In order to have an acceptable transport error in the estimates of $\mathring{R}_{q+1}$, the perturbation $w_{q+1}$ needs to be transported by the flow of the vector field $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$. Similar to [BV19], a natural way to achieve this, is to replace the linear phase $\xi\cdot x$ in the definition of the Beltrami wave $W_{\xi,\lambda}$ by the nonlinear phase $\xi\cdot\Phi(x,t)$, where $\Phi$ is transported by the aforementioned vector field. A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole line w.r.t time variable. For any integer $k\in\mathbb{Z}$, we subdivide $[k,k+1]$ into time intervals of size $\ell$ and solve transport equations on these intervals. For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$ we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
@@
We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follows from (C.5) and (C.6)
@@
As indicated by the proof in Appendix C, the validity of (4.8a) requires that both $v_{q}$ and $z_{q}$ have pointwise bounds, and this is why we need to cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b]. The above estimates play a crucial role in the cancellation of the estimates for the new Reynolds stress $\mathring{R}_{q+1}$. In particular, (4.8b) ensures the validity of stationary phase Lemma 4.2 below in Section 4.6, which requires the positive lower bounds of $\nabla\Phi_{k,j}$. We could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$. The extension is not unique but in the following we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$. We also let $\eta$ be a non-negative function supported in $(-1,1)$, which is 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
@@
form a partition of unity
$\sum_{j}\eta^{2}_{j}(t)=1,$
@@ 4.3.2. Construction of $w_{q+1}$
and
$\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell},$
where $c_{*}>0$ is a constant introduced in Lemma A.2 and we take the constant $c_{R}$ in (3.8f) satisfying $c_{R}<c_{*}$. Now, we define the amplitude functions
@@
$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}(t,x)}{\rho}\right)\right\|_{C^{0}}\leqslant c_{*}.$
@@
and
$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (4.15)
With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19], more precisely, for $t\in[k,k+1]$, we define
@@
where $B_{\xi}$ is suitable vector defined in Appendix A. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. Next, we define the incompressibility correction $w_{q+1}^{(c)}$. We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence free. To this end, it is useful to introduce the following scalar $phase$ function
@@
$W_{(\xi)}=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
Since from Appendix A $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$, and $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, we have
$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$
@@
$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k.k+1].$
@@
$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x)$
for all $t\in[k,k+1]$, $k\in\mathbb{Z}$. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted, and the new perturbation is defined as
$w_{q+1}=w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right),$ (4.19)
and so clearly $w_{q+1}$ is mean zero, divergence free and $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. We define the velocity field at level $q+1$ as
@@
Thus, it is also $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted.

### v3

@@ 4.1. Choice of parameters
In the sequel, additional parameters will be carefully chosen to verify all the conditions appearing in the estimates below. First, for a fixed integer $b\geqslant 7$, we let $\beta,\gamma\in(0,1)$ be small parameters satisfying
@@
which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$. The last free parameter $a\in 2^{36b^{2}\mathbb{N}}$ is chosen sufficiently large such that $a^{\frac{b\gamma}{8}}\in\mathbb{N}$ and $2<a^{(\frac{b}{4}-1)b^{2}\gamma}$. This ensures $f(q)\in\mathbb{N}$ and (3.6) holds. In the following sections, we increase $a$ to absorb various implicit and universal constants.
@@ 4.2. Mollification
In order to guarantee smoothness throughout the construction, we replace $v_{q}$ with a mollified velocity field $v_{\ell}$. For this purpose, we choose a small parameter $\ell=\lambda_{q}^{-2}$. Let $\{\phi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers on $\mathbb{R}^{3}$, and $\{\varphi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers with support in $(0,1)$. Here we use the one-sided mollifiers to reserve adaptedness. We define a mollification of $v_{q}$, $z_{q}$ and $\mathring{R}_{q}$ in space and time by
$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
where $\phi_{\ell}=\frac{1}{\ell^{3}}\phi(\frac{\cdot}{\ell})$ and $\varphi_{\ell}=\frac{1}{\ell}\varphi(\frac{\cdot}{\ell})$. By definition, it is easy to see that $z_{\ell}$, $v_{\ell}$ and $\mathring{R}_{\ell}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ adapted. Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ obey
$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{commutator}}),$ (4.1)
@@
where the commutator error $R_{\textrm{commutator}}$ and pressure term $p_{\ell}$ are given by
$R_{\textrm{commutator}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
@@ 4.3. Perturbation of the velocity
As usual in convex integration schemes, the new velocity field $v_{q+1}$ is a perturbation of $v_{\ell}$:
$v_{q+1}:=v_{\ell}+w_{q+1},$
where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed with construction by introducing flow maps that satisfy transport equations and time cutoffs.
@@ 4.3.1. Flow maps and cutoffs
In order to have an acceptable transport error in the estimates of $\mathring{R}_{q+1}$, the perturbation $w_{q+1}$ needs to be transported by the flow of the vector field $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$. Similar to [BV19], a natural way to achieve this, is to replace the linear phase $\xi\cdot x$ in the definition of the Beltrami wave $W_{\xi,\lambda}$ by the nonlinear phase $\xi\cdot\Phi(x,t)$, where $\Phi$ is transported by the aforementioned vector field. A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole time line ${\mathbb{R}}$.
For any integer $k\in\mathbb{Z}$, we subdivide $[k,k+1]$ into time intervals of size $\ell$ and solve transport equations on these intervals. For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$, we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
@@
We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follow from (C.5) and (C.6)
@@
As indicated by the proof in Appendix C, the validity of (4.4a) requires that both $v_{q}$ and $z_{q}$ have pathwise bounds. Therefore, we cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b]. The above estimates play a crucial role in canceling the oscillation error in the new Reynolds stress $\mathring{R}_{q+1}$. In particular, (4.4b) ensures the validity of stationary phase Lemma 5.3 below in Section 5.2, which requires the positive lower bounds of $\nabla\Phi_{k,j}$. Moreover, we could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$. The extension is not unique but in the following, we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$.
We also let $\eta$ be a non-negative function supported in $(-1,1)$, which equals to 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
@@
form a partition of unity, namely, for all $t\in[0,1]$
$\sum_{j}\eta^{2}_{j}(t)=1.$
@@ 4.3.2. Construction of $w_{q+1}$
and $\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell}$. The constant $c_{*}>0$ is introduced in Lemma A.2 and we take the constant $c_{R}$ in Proposition 3.3 satisfying
$c_{R}<c_{*}.$ (4.8)
Now, we define the amplitude functions
@@
$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\leqslant c_{*}.$
@@
With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19]. More precisely, for $t\in[k,k+1]$, we define
@@
where $B_{\xi}$ is a suitable vector defined in Appendix A. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted, we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted.
Next, we define the incompressibility correction $w_{q+1}^{(c)}$. We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence-free. To this end, it is useful to introduce the following scalar $phase$ function
@@
$W_{(\xi)}(x)=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
Since $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, it follows from Appendix A that $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$. By a direct computation, we deduce
$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$ (4.13)
@@
$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k,k+1].$
@@
$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x).$
Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}$ are $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted, we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted. Moreover, by (4.13) we deduce that
$w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right).$
Finally, the new perturbation is defined as
$w_{q+1}:=w^{(p)}_{q+1}+w^{(c)}_{q+1},$ (4.15)
and so clearly $w_{q+1}$ is mean zero, divergence-free and $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted. The new velocity at the level $q+1$ is defined as
@@
Thus, by the previous discussion, it is also divergence-free and $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted.
4.4. Definition of the Reynolds stress $\mathring{R}_{q+1}$
Recalling the system (4.1)
$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{commutator}}),$
[…]
$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.16)
$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
[…]
$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
In order to define the remaining oscillation error from the second line in (4.16), we first note that for $j$ and $j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$. Therefore, we may use Lemma A.2 and (4.10) to obtain
Note that compared to [BV19], a slight difference is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$. This term can be put into the pressure term $p_{\textrm{oscillation}}$. Above, we have denoted
and
Finally we define the Reynolds stress at the level $q+1$ by
and
5. Inductive estimates and proof of Proposition 3.3
In this section, we collect all the necessary estimates of $v_{q+1}$, $\mathring{R}_{q+1}$ and the energy to complete the proof of the Proposition 3.3. As a preliminary step, we need to give estimates on the $C_{t,x}^{N}$-norm of the amplitude functions $a_{(\xi)}$ defined in Section 4.3. This is done in forthcoming Proposition 5.1, whose proof is given in Appendix B.
Proposition 5.1.
Let $a_{(\xi)}$ be defined by (4.9). Then we have for any $N\geqslant 2$
and
$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma},$ (5.2)
where the implicit constants are independent of $q$.
This section is organized as follows: Section 5.1 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates. We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 5.2. Finally, in Section 5.3, we show how the energy is controlled. Note that we first give pathwise estimates by using (5.1) and (5.2), then taking expectation and using inductive assumptions can derive uniform moment estimates independent of time. In addition, we also need the estimates of transport equations, as provided in the appendix C. A useful stationary phase lemma on estimating $\mathring{R}_{q+1}$ previously used in the deterministic setting is recalled in Section 5.2.